OpenStudy (anonymous):

A hydrogen atom has a mass of 1.67kg. Find the mass of 71 x10^3 hydrogen atoms. Express the result in scientific notation

OpenStudy (anonymous):

hydrogen atom = 1.67 kg????? dont think so

OpenStudy (anonymous):

nevertheless its 1.67 * 71 * 10^3 mass = 118.57 * 10^3 = 1.1857 x 10^5

OpenStudy (anonymous):

note scientific notation = n.nnn.. x 10^n